How do I solve this equation? The directions tell me to "write the equation in standard form with integer coefficients"
1) 4x-y-7=0
2) -4x+5y+16=0
3) x-5=0
4) y=-5x+2
5) y=9x+1/2
This is algebra 1.
Thanks a lot!

Answer by ankor@dixie-net.com(22740)

"write the equation in standard form with integer coefficients"
You are not solving this, just putting the equation in the standard form which is:
Ax + By = C, where A, B, and C are integers
:
1) 4x-y-7=0
Add 7 to both sides
4x - y = 7
:
2) -4x+5y+16=0
Subtract 16 from both sides
-4x + 5y = -16
:
3) x-5=0
add 5 to both sides
x = 5; y coefficient is 0
:
4) y=-5x+2
add 5x to both sides
5x + y = 2
:
5) y=9x+1/2
subtract 9x from both sides
-9x + y = 1/2
multiply by 2 to get rid of the fraction
-18x + 2y = 1